OpenStudy (precal):

integral -2 to 3 of ((1/x^2) +x)dx

OpenStudy (precal):

\[\int\limits_{-2}^{3}\left( \frac{ 1 }{ x^2}+x \right)dx\]

OpenStudy (precal):

I might have posted this one before but I don't recall. Calculator won't evaluate this integral. I doublechecked the graph, is it because 0 is included

OpenStudy (precal):

and 0 is not defined in the domain

OpenStudy (anonymous):

you have a problem at 0. break it up into two intervals... -2 to 0 and 0 to 3. you'll have to take the limits at 0.

OpenStudy (anonymous):

you gotta do the improper integral

OpenStudy (kainui):

You're using the wrong calculator, try this one: http://www.wolframalpha.com/input/?i=integral%20from%20-2%20to%203%20(x%5E-2%2Bx)dx&t=crmtb01

OpenStudy (anonymous):

\[\int\limits_{-2}^{3}\left( \frac{ 1 }{ x^2 }+x \right)dx=\lim_{t \rightarrow 0}\left[ \int\limits_{-2}^{t}\left( \frac{ 1 }{ x^2 }+x \right)dx+\int\limits_{t}^{3}\left( \frac{ 1 }{ x^2 }+x \right)dx \right]\]

OpenStudy (anonymous):

make sense?

OpenStudy (precal):

yes, but I think this problem is in the incorrect place. I believe improper integrals is not in my course but the next course. Thanks for the set up

OpenStudy (anonymous):

first sem calc? should be in there

OpenStudy (turingtest):

I didn't see em till calc II

OpenStudy (precal):

not label as improper integral or is this not improper integral?

OpenStudy (precal):

I am interested in seeing how to finish solving it using the limit approach. If nothing else, just for fun :)

OpenStudy (turingtest):

In calculus, an improper integral is the limit of a definite integral as an endpoint of the interval(s) of integration approaches either a specified real number or ∞ or −∞ or, in some cases, as both endpoints approach limits - Wikipedia so technically it's not an improper integral as the problem is presented, but if you want to show that it diverges, you need to make it into two improper integrals

OpenStudy (anonymous):

you do as normal but you take the limit when evaluating at the trouble point

OpenStudy (precal):

see I didn't see any diverges or converges problems until Calculus III

OpenStudy (anonymous):

i find this site much better for this level... http://tutorial.math.lamar.edu/Classes/CalcII/ImproperIntegrals.aspx

OpenStudy (turingtest):

I just wanted to explain why it wasn't improper as it was originally stated, but we made it improper

OpenStudy (precal):

np thanks

OpenStudy (anonymous):

\[\lim_{t \rightarrow 0}\int\limits_{-2}^{t}\left( \frac{ 1 }{ x^2 }+x \right)dx=\lim_{t \rightarrow 0}\left( -\frac{ 1 }{ x } +\frac{ x^2 }{ 2 }\right)_{-2}^{t}=\]

OpenStudy (anonymous):

check out Paul's online notes (previously posted). really quite good.

OpenStudy (precal):

thanks again.

OpenStudy (anonymous):

you're welcome! btw, i only did part of the integral above, but you probably realized that!

OpenStudy (precal):

that is ok, gotta run an errand. Be back in a little bit....sorry I can't stay